OpenStudy (anonymous):

It takes 1629 digits to number the pages of a book. How many pages does the book have?

OpenStudy (anonymous):

1629/2

OpenStudy (anonymous):

846 pages

OpenStudy (anonymous):

I think it shud be 579. I got this, can someine confirm.

OpenStudy (anonymous):

how did you get 579?

OpenStudy (anonymous):

From page #1 to page # 9, there are 9 pages From page #10 to page # 99, there are 90 pages. However, since each page number has 2 digits, there are 180 digits from page # 10 to page # 99 From page #100 to page # 999, there are 900 pages. However, since each page number has 3 digits, there are 2700 digits from page # 100 to page # 999 Since there are 1629 digits, we know so far that number of pages is less than 900 Now, we need to find out how many digits there are for those pages greater than 99 or those pages with 3 digits Since there are 189 digits for pages with less than 3 digits, we can subtract 189 from 1629 digits to find the number of pages with 3 digits 1629 − 189 = 1440 1440 represents the pages with 3 digits, so number of pages with 3 digits = 1440/3 = 480 Thus, total number of pages = 480 + 90 + 9 = 579 pages

OpenStudy (anonymous):

um... i think that you are over-thinking this....

OpenStudy (anonymous):

it would just be divided by 2... right?

OpenStudy (anonymous):

hes correct

OpenStudy (anonymous):

who is correct?

OpenStudy (anonymous):

Adil.

OpenStudy (karatechopper):

joe did u teach karate?

OpenStudy (anonymous):

ok, well then i guess I was wrong

OpenStudy (anonymous):

or at least the reasoning is, im going to try and verify.

OpenStudy (karatechopper):

or self defense?

OpenStudy (anonymous):

You don't number the first page(s) of a book... perhaps it's a 580 ;P

OpenStudy (karatechopper):

joe?

OpenStudy (across):

AdilAshraf is correct.

OpenStudy (anonymous):

Basically the question is asking, if you write down the natural numbers as such: 1234567891011121314151617... and you go to the 1629th digit, what number do you end on?

OpenStudy (karatechopper):

joe could u please go to chat im math?

OpenStudy (anonymous):

i dont teach karate unfortunately =/

OpenStudy (karatechopper):

well u look like my teacher... for self defense. did u teach that?

OpenStudy (anonymous):

not at all =/

OpenStudy (karatechopper):

hmm do you you have a twin?

OpenStudy (anonymous):

You refering to the 'self defense wis jo' ? Those youtubes are awesome xD

OpenStudy (anonymous):

Nope no twin that i know of.

OpenStudy (anonymous):

lol "that I know of"

OpenStudy (karatechopper):

wow. my teacher was named joe too and ur pic looks just like him!

OpenStudy (anonymous):

well there are always people that look like other people out there. I dont want to claim "im the only person that looks like me ever!" lol

OpenStudy (karatechopper):

wow this is justt crazy